OpenStudy (anonymous):

Solve for x. 8^(9x)=3^(x-9) Write the exact answer using base-10 logarithms.

OpenStudy (anonymous):

Take Log both the sides, \[\huge 9xLog_{10}8 = (x-9)Log_{10}3\]

OpenStudy (anonymous):

\[\implies 9xLog_{10}2^3 = (x-9)Log_{10}3\] \[\implies (9x \times 3) Log_{10}2 = (x-9)Log_{10}3\]